### v1

# Simplifying Triangle Meshes in the Wild

Abstract.
This paper introduces a fast and robust method for simplifying surface triangle meshes in the wild while maintaining high visual quality. While previous methods achieve excellent results on manifold meshes by using the quadric error metric, they suffer from producing high-quality outputs for user-created meshes, which often contain non-manifold elements and multiple connected components. In this work, we begin by outlining the pitfalls of existing mesh simplification techniques and highlighting the discrepancy in their formulations with existing mesh data. We then propose a method for simplifying these (non-manifold) triangle meshes, while maintaining quality comparable to the existing methods for manifold inputs. Our key idea is to reformulate mesh simplification as a problem of decimating simplicial 2-complexes. This involves a novel construction to turn a triangle soup into a simplicial 2-complex, followed by iteratively collapsing 1-simplices (vertex pairs) with our modified quadric error metric tailored for topology changes. Besides, we also tackle textured mesh simplification. Instead of following existing strategies to preserve mesh UVs, we propose a novel perspective that only focuses on preserving texture colors defined on the surface, regardless of the layout in the texture UV space. This leads to a more robust method for textured mesh simplification that is free from the texture bleeding artifact. Our mesh simplification enables level-of-detail algorithms to operate on arbitrary triangle meshes in the wild. We demonstrate improvements over prior techniques through extensive qualitative and quantitative evaluations, along with user studies.
journal: TOG
1. Introduction
Mesh simplification creates different levels of detail (LOD) of 3D objects while maintaining their visual fidelity. It plays a critical role in interactive graphics to achieve the required performance on rendering and simulation across various hardware. Its importance has been motivating decades of development with a plethora of solutions, such as the widely used QEM by Garland and Heckbert [1997]. Though these solutions can produce high-quality results when the input is a manifold surface, we demonstrate that they fail to generate acceptable results on a majority of meshes obtained from online repositories (see Figure 1, Figure 2).
Figure 2. Off-the-shelf mesh simplification techniques (e.g., [Garland and Heckbert, 1998]) produce excellent results when the input mesh meets the common assumption of being a single manifold mesh (top green). However, when the input mesh does not meet the input assumption, existing methods often fail to achieve satisfactory results (bottom green). As these “defective” inputs form the majority of human-created meshes, this motivates our work: a mesh simplification method that can handle both cases (blue). Note that for the implementation of [Garland and Heckbert, 1998], we use the version from [Cignoni et al., 2008] with topology preservation.
This issue is caused by the discrepancy between existing problem formulations and real-world situations. Existing LOD techniques often assume the input mesh is manifold [Dey et al., 1999]. This was a realistic assumption when, e.g. in the 90s, the goal was to simplify meshes extracted from isosurfacing methods, such as marching cubes [Lorensen and Cline, 1987]. However, recent advancements in 3D modeling have populated online repositories with a rich amount of manually modeled shapes. In contrast to meshes reconstructed from isosurfacing, a majority of them are non-manifold (e.g., 98.9% on ShapeNet [Chang et al., 2015] are non-manifold, see Figure 3). The significant amount of non-manifold meshes demonstrates the importance of revisiting mesh simplification and its problem formulation based on today’s scenario.
Figure 3. Many meshes one can encounter in online repositories are non-manifolds. For instance, Thingi10k [Zhou and Jacobson, 2016] (a relatively clean data with only 3D-printable shapes) possesses 22.5% non-manifold shapes, ModelNets [Wu et al., 2015] (a dataset of CAD models) contains 54.3% non-manifold objects, and more severely 98.9% of the meshes in the ShapeNet dataset [Chang et al., 2015] (a widely used dataset for machine learning) are non-manifolds.
A desired mesh simplification method must be performant and robust to all types of geometric artifacts. In real-time applications such as virtual reality and video games, these (defective) meshes appear frequently during game time from user interactions (i.e., engineers collaboratively model 3D buildings). Being able to efficiently and robustly simplify all the newly created 3D content on the fly is critical to achieving target performance. Such high performance requirements also exclude expensive alternatives, such as repairing defective meshes into manifolds and performing simplification afterward, which may take minutes to hours.
Figure 4. We stress test our method by simplifying a variety of challenging cases encountered in the wild, including noisy 3D scans (left), low-quality triangles from CAD models (middle), and meshes with defects (right).
In this work, we propose a robust method to efficiently simplify triangle meshes that may contain multiple components, non-manifolds, and even the extreme case like triangle soups (Figure 4). Unlike previous methods that assume manifold inputs, we elevate the restriction and allow our approach to work with a soup of triangles. This enables us to handle all triangle meshes one can encounter today. To handle textured meshes, instead of preserving UVs like prior arts, we propose to preserve textures during simplification via successive mapping.
This paper starts with a detailed analysis of the pitfalls of existing mesh simplification strategies and a combination of them (section 4). We then detail our robust mesh simplification method which consists of (1) a construction strategy of simplicial 2-complex tailored for mesh simplification (section 5), (2) a modified quadric error metric with a small change in the metric accumulation strategy that leads to significant improvements in the output quality (section 6), and (3) a novel texture preservation simplification strategy based on successive mapping (section 7). We demonstrate that our method can robustly simplify triangle meshes in the wild and achieve better results than prior arts on qualitative, quantitative, and user study evaluations (section 9). Our work contributes to the emerging toolkit for robust geometry processing, such as [Jacobson et al., 2013; Sawhney and Crane, 2020], by offering a robust mesh LOD technique.
2. Related Work
Our method belongs to the family of local decimation schemes defined on triangle meshes, aiming to preserve their geometric appearance. We thus focus our discussion on prior techniques closest to ours, and only briefly touch other formulations, such as global optimization and computation-preserving simplification. We omit the discussion on simplifying other shape representations, such as quadrilateral meshes, implicit functions, and point clouds, and refer readers to Luebke [2003] for a more comprehensive discussion.
Triangle Mesh Simplification
Mesh simplification has been extensively studied in computer graphics to reduce the resolution of 3D objects while preserving their appearance. Early attempts try to fit low resolution polygons [Jr. and Zyda, 1991], globally remesh the input to a lower resolution [Hoppe et al., 1992; Turk, 1992], or locally remove each mesh element [Schroeder et al., 1992]. Since then, different formulations have been proposed, such as clustering of triangles [Cohen-Steiner et al., 2004; Xu et al., 2024] or vertices [Rossignac and Borrel, 1993; Low and Tan, 1997; Lindstrom, 2000], greedily removing triangles [Hamann, 1994; Gieng et al., 1998] or vertices [Klein, 1998], simplifying a manifold envelope wrapping around the input [Mehra et al., 2009; Gao et al., 2022; Chen et al., 2023b], or optimization with a differentiable renderer [Hasselgren et al., 2021; Deliot et al., 2024; Knodt et al., 2024]. Some of these techniques have also been revisited with learning-based techniques [Potamias et al., 2022; Chen et al., 2023a]. Among them, one of the most popular strategies is the iterative edge collapse [Hoppe, 1996] with the quadric error metric [Garland and Heckbert, 1997]. The popularity of edge collapses is partly due to the high-quality output from the quadric error metric and its efficiency, exhibiting constant time complexity for each edge collapse.
Quadric Error Metrics
The success of quadric error [Garland and Heckbert, 1997] has been motivating the development of a variety of quadrics for measuring different geometric quantities. Lindstrom and Turk [1998] propose volume and area quadrics to measure volume and area changes, respectively. Garland and Heckbert [1998] introduce the boundary quadric to maintain shape boundaries. Trettner and Kobbelt [2020] consider a mesh element as a sample from a probability distribution and derive the probabilistic quadric to further improve its numerical behavior. Other extensions generalize the quadric error to incorporate surface attributes (e.g., textures) by defining quadric error in a higher dimension [Garland and Heckbert, 1998; Garland and Zhou, 2005] or by adding extra quadric terms derived from a linear approximation of the attribute [Hoppe, 1999]. The optimal choice of metric varies depending on the input geometry and applications. The suite of quadric error metrics provides us with several options to mix-and-match them for different tasks, including the ones beyond mesh simplification such as quadric-inspired surface reconstruction [Kobbelt et al., 2001; Ju et al., 2002; Schaefer and Warren, 2005] and filtering [Vieira et al., 2010; Legrand et al., 2019].
Topology Varying Mesh Simplification
Many of the methods mentioned above are designed for simplifying a single manifold triangle mesh. Another class of mesh simplification generalizes the local decimation method to triangle meshes in the wild, which may contain multiple components or non-manifolds. Garland and Heckbert [1997] reformulate the edge collapse operation as an operation to collapse any vertex pair, meaning one can merge two vertices even if they are not connected by an actual edge. These vertex pairs are computed based on vertex-to-vertex distance (see Figure 12) and optimized with the quadric error metric. Instead of vertex pair collapses, Schroeder [1997] identifies different cases for manifold/non-manifold vertices and removes each vertex based on a distance-to-plane metric. If camera information is given, El-Sana and Varshney [1999] demonstrate that one can incorporate camera information into the simplification, In order to handle any triangle meshes, Popovic and Hoppe [1997] reformulate it as a problem of simplifying simplicial complex. Our method is another instance of topology varying simplification for simplicial complex with high quality geometric approximation and is capable of transferring textures.
Another possible direction is to repair the mesh, i.e. convert it into a single manifold, and then simplify it. However, this depends on a repairing method to produce high quality outputs, which is not always the case (see Figure 5).
Figure 5. Repairing a triangle mesh with [Hu et al., 2020] (third) and then simplifying the repaired model (fourth) depends on a having high quality repaired model. However, this is not always the case. Our method simplifies the input mesh directly and leads to a better result (first).
Alternatively, one can robustly create a manifold “cage” to wrap the triangle mesh and then simplify the manifold wrapper [Nooruddin and Turk, 2003; Chen et al., 2023b]. However, these methods, e.g. Chen et al. [2023b], often suffer from deleting interior components which may be important for interacting with the 3D object (see Figure 6).
Figure 6. Compared to methods that construct a manifold mesh wrapper and then simplify it [Chen et al., 2023b] (middle), our method directly simplifies the input and preserves interior structures (right).
We thus follow the spirit of early attempts to simplify triangle meshes with vertex pair collapses.
Beyond Appearance Preservation
Mesh simplification has also been studied in the context of preserving simulation-related properties, such as preserving acoustic transfer [Li et al., 2015], spectral properties [Liu et al., 2019; Chen et al., 2020; Keros and Subr, 2023; Lescoat et al., 2020], intrinsic geometry [Liu et al., 2023; Shoemaker et al., 2023], and the dynamic behavior of physics-based simulations [Kharevych et al., 2009; Chen et al., 2015, 2017, 2018]. When trained jointly with a machine learning model, one could even tailor a simplification for the underlying task, such as mesh classification [Hanocka et al., 2019; Ludwig et al., 2023]. These methods serve as fundamental building blocks to obtain a coarse approximation for a simulation. But if one aims to build a multilevel solver that requires to transfer signals across resolutions [Guskov et al., 1999], an important ingredient is to compute correspondences across the decimation hierarchy. This motivation stimulated several works on computing correspondences across hierarchy by an intrinsic mapping for each local decimation [Lee et al., 1998; Khodakovsky et al., 2003; Liu et al., 2020, 2021, 2023] or by maintaining a bijective map with extrinsic projection [Jiang et al., 2020, 2021]. With such correspondences, one can then build hierarchical solvers for scalable computation [Liu et al., 2021; Zhang et al., 2022; Wiersma et al., 2023; Zhang et al., 2023].
Our method for texture transfer is an instance of simplification that maintains correspondences. But the key difference is that our input may contain non-manifold structures and our method is designed to introduce topological changes. These cases prevent us from leveraging existing mesh parameterization to ensure e.g. bijectivity with low distortion. We thus fall back to a progressive closest point projection, as detailed in section 7, for transferring textures.
3. Background
Our method simplifies a triangle mesh by iterative edge collapses. We prioritize the sequence of edge collapses by our modified quadric error metric. Here, we briefly introduce the quadric error metric [Garland and Heckbert, 1997] and some variants that are relevant to our method.
3.1. Quadric Error Metric
Let ${\mathbf{p}}\in\mathcal{M}$ be a point on a surface $\mathcal{M}$ embedded in $\mathbb{R}^{3}$ and $\hat{{\mathbf{n}}}$ be the (unit) normal vector at ${\mathbf{p}}$. The tangent plane at ${\mathbf{p}}$ is given by all points ${\mathbf{x}}\in\mathbb{R}^{3}$ that satisfy
(1) $\hat{{\mathbf{n}}}^{\top}({\mathbf{x}}-{\mathbf{p}})=0.$
The quadric error $E({\mathbf{x}})$ measures the squared distance from any point ${\mathbf{x}}$ to the tangent plane at ${\mathbf{p}}$, which can be computed as
(2) $E({\mathbf{x}})$ $=\big{(}\hat{{\mathbf{n}}}^{\top}({\mathbf{x}}-{\mathbf{p}})\big{)}^{2}$
(3) $=({\mathbf{x}}-{\mathbf{p}})^{\top}\hat{{\mathbf{n}}}\hat{{\mathbf{n}}}^{\top}({\mathbf{x}}-{\mathbf{p}})$
(4) $={\mathbf{x}}^{\top}{\mathbf{A}}{\mathbf{x}}+2{\mathbf{b}}^{\top}{\mathbf{x}}+c,$
with
(5) ${\mathbf{A}}=\hat{{\mathbf{n}}}\hat{{\mathbf{n}}}^{\top},\quad{\mathbf{b}}=-{\mathbf{A}}{\mathbf{p}},\quad c={\mathbf{p}}^{\top}{\mathbf{A}}{\mathbf{p}}$
A quadric $Q$ refers to the triplet $({\mathbf{A}},{\mathbf{b}},c)$, which are quantities derived from the plane equation $\mathcal{P}$ with a unit normal $\hat{{\mathbf{n}}}$ and a point ${\mathbf{p}}$ on the tangent plane:
(6) $Q\coloneq({\mathbf{A}},{\mathbf{b}},c)=\mathcal{P}(\hat{{\mathbf{n}}},{\mathbf{p}}),$
This quadric $Q$ gives us the complete information to compute the quadric error $E({\mathbf{x}})$.
Triangle Quadrics
The quadric error can be generalized to triangulated 2-manifolds by defining the quadric error $E_{ijk}$ on the plane of each triangle $ijk$
(7) $E_{ijk}({\mathbf{x}})={\mathbf{x}}^{\top}{\mathbf{A}}_{ijk}{\mathbf{x}}+2{\mathbf{b}}_{ijk}^{\top}{\mathbf{x}}+c_{ijk}.$
This defines the triangle quadric $Q_{ijk}$ as
(8) $Q_{ijk}=({\mathbf{A}}_{ijk},{\mathbf{b}}_{ijk},c_{ijk})=\mathcal{P}({\mathbf{n}}_{ijk},{\mathbf{v}}_{i})$
derived from the face normal $\hat{{\mathbf{n}}}_{ijk}$ and the location of a point on the plane, such as one of the triangle corner vertices ${\mathbf{v}}_{i}$.
Vertex Quadrics
For each vertex $i$, the vertex quadric error $E_{i}$ is defined as the weighted summation of the triangle quadric errors $E_{ijk}$ from its one-ring triangles $ijk$
(9) $E_{i}({\mathbf{x}})$ $=\sum_{ijk\in\mathcal{N}_{i}}a^{i}_{ijk}E_{ijk}({\mathbf{x}})$
(10) $=\sum_{ijk\in\mathcal{N}_{i}}a^{i}_{ijk}({\mathbf{x}}^{\top}{\mathbf{A}}_{ijk}{\mathbf{x}}+2{\mathbf{b}}_{ijk}^{\top}{\mathbf{x}}+c_{ijk})$
[…]
We use $\mathcal{N}_{i}$ to denote the one-ring triangles of the vertex $i$ and $a^{i}_{ijk}$ is a portion of the vertex area at $i$ coming from face $ijk$. A common choice of area is the barycentric area (see inset) which simply sets $a^{i}_{ijk}=a_{ijk}/3$ to be one-third of the face area $a_{ijk}$ [Meyer et al., 2002]. This derivation gives rise to the definition of vertex quadric $Q_{i}$ as a weighted summation of its one-ring triangle quadrics $Q_{ijk}$
(11) $Q_{i}=\sum_{ijk\in\mathcal{N}_{i}}\!\!a^{i}_{ijk}Q_{ijk}$
where the summation between quadrics is the component-wise summation for all elements in the triplets.
Edge Quadrics
The quadric error metric $E_{ij}$ for each edge $ij$ is defined as the summation of vertex quadric errors of its endpoints
(12) $E_{ij}({\mathbf{x}})=E_{i}({\mathbf{x}})+E_{j}({\mathbf{x}})$
As we can see in Equation 9 that summing up quadric errors leads to a summation of quadrics, this leads to the definition of an edge quadric $Q_{ij}$ as
(13) $Q_{ij}=Q_{i}+Q_{j}.$
3.2. Quadric Error Edge Collapses
Garland and Heckbert [1997] suggest simplifying a triangle mesh by iteratively collapsing the edge with the smallest edge quadric error $E_{ij}({\mathbf{x}}^{\star})$ which is the value of $E_{ij}$ evaluated at the optimal location ${\mathbf{x}}^{\star}$ that minimizes $E_{ij}$. Let $Q_{ij}=({\mathbf{A}}_{ij},{\mathbf{b}}_{ij},c_{ij})$, the optimal location ${\mathbf{x}}^{\star}$ can be computed solving a linear system obtained from setting $\nabla E_{ij}=0$
(14) ${\mathbf{A}}_{ij}{\mathbf{x}}^{\star}=-{\mathbf{b}}_{ij}$
If the matrix ${\mathbf{A}}_{ij}$ is invertible and well-conditioned, one can solve for ${\mathbf{x}}^{\star}$ with standard linear solvers, such as Cholesky decomposition. If not, one can use the singular value decomposition to solve for ${\mathbf{x}}^{\star}$ more robustly [Lindstrom, 2000] or use the probabilistic quadric as a regularization [Trettner and Kobbelt, 2020].
In subsequent iterations, instead of recomputing the edge quadric with Equation 13, Garland and Heckbert [1997] recommend using the $Q_{ij}$ as the vertex quadric for the newly inserted vertex. Suppose we collapse an edge $ij$ to a vertex $i^{\prime}$, the quadric of the new vertex $Q_{i^{\prime}}=Q_{ij}$ is simply the edge quadric before the collapse, and $Q_{i^{\prime}}$ will then be used to update the quadrics for its one-ring edges. This definition allows the quadric error to “memorize” all the plane information on the input mesh, instead of from the current mesh. In practice, compared to recomputing the edge quadric $Q_{ij}$ from scratch, such an accumulation often leads to a more efficient and desirable decimation. But some applications may favor not accumulating quadrics [Hoppe, 1999].
We use the term “memory” to denote the implementation that accumulates the quadric error metric throughout the decimation and “memoryless” to denote the implementation that recomputes the quadric at every iteration.
3.3. Area Quadrics
The quadric error metric has stimulated several variants to serve different purposes (see section 2). Among them, the area quadric is related to our decimation metric. The area quadric measures the squared area change for each edge collapse. Lindstrom and Turk [1998] show that the squared area change for each edge $ij$ can be written as a quadratic expression, thus giving birth to the area quadric $Q^{A}_{ij}$
(15) $Q^{A}_{ij}=\sum_{ab\in\mathcal{N}_{ij}\cap\partial\mathcal{M}}\!\!\Big{(}[{\mathbf{s}}_{ab}]_{\times}^{\top}[{\mathbf{s}}_{ab}]_{\times},\ -[{\mathbf{s}}_{ab}]_{\times}{\mathbf{t}}_{ab},\ {\mathbf{t}}_{ab}^{\top}{\mathbf{t}}_{ab}\Big{)}/2$
with
(16) ${\mathbf{s}}_{ab}={\mathbf{v}}_{b}-{\mathbf{v}}_{a}\qquad{\mathbf{t}}_{ab}={\mathbf{v}}_{a}\times{\mathbf{v}}_{b}$
where $ab$ are $ij$’s one-ring edges that are also on the mesh boundary $\partial\mathcal{M}$ and $[\cdot]_{\times}$ is the cross product matrix defined as
(17) $[{\mathbf{x}}]_{\times}\coloneq\begin{bmatrix}0&-x_{2}&x_{1}\\ x_{2}&0&-x_{0}\\ -x_{1}&x_{0}&0\end{bmatrix}$
This area quadric $Q^{A}_{ij}$ is derived from summing up the squared area (computed with cross products) from each boundary edge $ab$ to the newly inserted vertex $i^{\prime}$ (see inset). This implies that this area quadric is merely an approximation to the actual area change because this measure ignores the cancelation between the area gain and the loss. But this proxy maintains the favorable quadratic expression and can be seamlessly incorporated into the quadric-based mesh simplification.
4. Pitfalls of Mesh Simplification in the Wild
This section analyzes some common artifacts when deploying off-the-shelf algorithms to simplify meshes in the wild. Off-the-shelf methods often assume manifold inputs with few (or even a single) connected components. Despite being a realistic assumption for meshes extracted from e.g. marching cubes, this assumption becomes obsolete when one considers meshes on online repositories, where many of them are created by artists. The discrepancy in the actual input and the input assumption leads to several geometric and texture artifacts. Even if one mixes-and-matches several existing ingredients, they still lead to undesired geometric and texture distortion. In practice, this implies that artists require to meticulously paying attention to the mesh connectivity in ways that favors existing LOD algorithms. This not only burdens the 3D modeling process and prevent democratizing content creation for inexperienced users. In this section, we highlight some pitfalls in existing mesh simplification techniques, motivating the need to revisit (textured) mesh simplification with a new problem formulation that is more aligned with existing 3D models.
4.1. Pitfalls in Geometric Error
Figure 7. Given a tree mesh with non-manifold elements and multi-connected components, we demonstrate that several off-the-shelf simplification methods and a mixture of them suffer from different issues, such as having a high resolution lower bound or deleting planar components. In contrast, our method leads to perceptually better results when decimating the mesh down to 1% of the input resolution, supported by the user study in subsection 9.3.
Undesired Topology Preservation
Several mesh simplification methods, such as [Dey et al., 1999], aim at maintaining the mesh topology during simplification. However, maintaining topology implicitly sets a lower bound on the coarsest mesh resolution that depends on the number of connected components. This is because one has to use at least a single triangle to represent a component. For instance, given a tree mesh, separate geometric parts are often modeled as different connected components (e.g., each leaf is a distinct component). Thus, using topology-preserving simplification quickly hits the lower bound and prevents further decimation (see Figure 7 A). This restriction becomes problematic for low-end devices (e.g., mobiles, headsets), which require aggressive mesh decimation to meet strict resolution limits. Not to mention that many input meshes in practice already possess topological artifacts to begin with, which will break implementations that assume manifold inputs immediately. The issue of having topological artifacts has been motivating the development of robust algorithms in different areas, such as rendering and collision detection, so that we can still “consume” these meshes in today’s e.g. video games. This further motivates us to remove the topology preservation in mesh simplification to accommodate the real-time applications we consider in our work.
Area Loss
If we elevate the topology preservation mentioned above and use a state-of-the-art quadric error metric by Trettner and Kobbelt [2020] for decimation, the mesh simplification still suffers from unexpected results (see Figure 7 B). This is because several quadric error metrics, such as [Trettner and Kobbelt, 2020], measure errors as the squared distance to planes. When a mesh component is closer to a plane (e.g., leaves) or a developable patch, quadric error will report a small error when collapsing edges at the planar/developable region, regardless of how it may impact the geometric error. This results in the behavior of removing large components even though such a removal may cause large geometric errors.
No Virtual Edge
A potential approach to address the area loss is to incorporate the area quadric introduced by Lindstrom and Turk [1998], in addition to the edge quadric error. In Figure 7 C, we demonstrate the result of combining the area quadric [Lindstrom and Turk, 1998] and the probabilistic quadric [Trettner and Kobbelt, 2020] (we use [L98+T20] as the abbreviation of this combination). Although it helps to preserve some large-area components, it is still limited to deleting components, instead of finding a better approximation by merging small components into bigger ones. The reason is that such a combination only has the degrees of freedom to collapse existing edges. These edge collapses will not merge disconnected parts because they do not share an edge.
Limitations of Vertex Virtual Edges
In order to encourage merging components (e.g., trunk and branches), constructing “virtual edges” across different components is desired. Garland and Heckbert [1997] suggest constructing virtual edges for vertex pairs such that their vertex-to-vertex distance is below a small threshold. One can incorporate such a strategy, a combination of [Trettner and Kobbelt, 2020; Lindstrom and Turk, 1998; Garland and Heckbert, 1997] (called [G97+L98+T20] for short), to decimate a 3D model and expect it to merge disconnected components. However, in Figure 7 D, we show that this strategy still does not introduce significant improvements. This is because the vertex distance is often an inadequate metric to measure the intended connectivity by artists. In the inset, we zoom in on the intersection between the trunk and a branch. We can notice that connectivity is implied by surface proximity, instead of making sure vertices are colocated. This results in failure in constructing virtual edges that can effectively merge components. One may go down to the direction of increasing the threshold of vertex distances and fine-tune the threshold. However, determining an appropriate threshold is not trivial because a large threshold results in virtual edges being connected excessively, and a desired threshold value often varies across different regions of the same mesh due to non-uniform tessellation. The limitation in the vertex distance measure also generalizes to several artist-created assets, such as the lamp model in Figure 12. Even if one can identify a great threshold for vertex distances, it still leads to over smoothed results detailed later in Figure 15.
4.2. Pitfalls in Texture Error
Besides geometric issues, off-the-shelf LOD tools also encounter issues in preserving surface attributes, such as textures.
UV coordinates are a popular way to store the mapping to a texture space that stores surface attributes. A traditional approach to preserve mesh attributes during simplification is to minimize geometric distortion of their UVs [Garland and Heckbert, 1998; Hoppe, 1999]. However, such a strategy of approximately preserving UV coordinates only works well when the input mesh has only few connected components in the UV domain, a.k.a. UV islands. This is because if the boundary of each UV island is not exactly preserved, and when the renderer interpolates texture colors near UV boundaries, distortion in the boundary curves could lead to interpolating the background color of the UV image on to the surface mesh. This behavior of rendering the background color on the mesh surface is also known as the “texture bleeding” artifact (see Figure 8).
Figure 8. Existing textured mesh simplification techniques work well when the input mesh has one or few connected components in the UV-space, a.k.a. UV islands (bottom middle). However, when deploying to meshes that have multiple UV islands, previous methods often lead to the “texture bleeding” artifact: seeing the background color of the texture image on the surface mesh (top middle). As most real-world meshes (e.g., [Maggiordomo et al., 2020]) contain multiple texture islands, this motivates our method to handle both cases more robustly (right column).
In practice, a majority of the meshes have more than one UV islands, leading to significant texture bleeding (see Figure 8) or texture distortions (see Figure 9) when applying this strategy to meshes in the wild. In addition, striving preservation in the UV coordinates could deteriorate the quality of geometry decimation (see Figure 19).
Figure 9. Previous methods simplify a mesh while preserving its UV-coordinates for transferring texture to the simplified mesh. In addition to texture bleeding artifacts, they could lead to distorted textures. Our method reduces such a distortion by computing correspondences between the input and its simplified counterpart (right).
On the other hand, even though methods like differentiable rendering [Hasselgren et al., 2021] could potentially repair texture bleeding, but such a texture optimization often leads to minutes to hours of additional computation time (see Figure 10), making them unsuitable for real-time applications mentioned in section 1,
Figure 10. Compared to an expensive mesh optimization method with differentiable rendering that usually takes a few hours [Hasselgren et al., 2021], our method is orders of magnitudes faster (see section 9) and achieves comparable quality.
An alternative is to maintain the boundary shape of each UV island exactly [Liu et al., 2017]. However, this can lead to significant geometric distortion and possibly result in the early termination of the simplification process (see Figure 11).
Figure 11. An alternative method to avoid texture bleeding is to preserve the UV boundaries exactly, such as [Liu et al., 2017]. But they often lead to larger geometric distortion and have a lower bound on the resolution of the simplified mesh (top rows). Our method, in contrast, can decimate the method aggressively while still being able to transfer texture colors.
4.3. Baselines
In this section, we highlight the pitfalls of existing mesh simplification techniques. We show that simply using techniques proposed in single papers (e.g., [Trettner and Kobbelt, 2020]) cannot achieve satisfying results on simplifying non-manifold and multi-component meshes. We further show that naive combinations of existing techniques, including [L98+T20] ([Lindstrom and Turk, 1998; Trettner and Kobbelt, 2020]) and [G97+L98+T20] ([Garland and Heckbert, 1997; Lindstrom and Turk, 1998; Trettner and Kobbelt, 2020]) suffer from different issues. Not to mention that these works solely focus on geometry simplification, they do not handle textures. Thus, in order to comprehensively evaluate our approach, we evaluate the geometric quality against papers that solely focus on geometry simplification, such as [Garland and Heckbert, 1997], [Trettner and Kobbelt, 2020], and our invented combinations ([L98+T20] , [G97+L98+T20] ). For textured mesh simplification, we evaluate our method against individual works presented in subsection 4.2. We will show that our method leads to improvements over both geometry and textured mesh simplification.
5. Virtual Edges & Edge Collapses
Given a triangle mesh with vertices $V$ and faces $F$, the first step of our method is to construct a set of (physical and virtual) edges $E$. We use physical edges to denote the vertex pairs that directed connected in $F$ and virtual edges to denote the pairs that do not exist in $F$ (see inset). The combination $\mathcal{M}=(V,E,F)$, a.k.a. simplicial 2-complex, is the input to our system. We then decimate the model by iteratively collapsing each edge in $E$ prioritized by our modified quadric error metric presented in section 6.
5.1. Virtual Edges with Triangle Distances
Defining virtual edges is crucial when decimating a scene with multiple components. Without them, a simplification algorithm would be limited to deleting each connected part, rather than merging them to potentially achieve a better approximation (see Figure 13).
Garland and Heckbert [1997] propose to construct a virtual edge if the Euclidean distance between two vertices is below a small threshold. Despite working well for triangle soups coming from dicing up a 2-manifold triangle mesh (e.g., the bottom row in Figure 14), this strategy often fails on meshes in the wild because it cannot reliably detect whether two components are connected. Vertices are merely a sparse set of samples from the underlying surface. The distance computed from them often incorrectly overestimate the true distance between two components. Meshes in the wild often contain visually connected pieces formed by topologically disconnected parts with no visible gap between them. This can lead to large vertex-to-vertex distance due to the topology of the mesh, e.g. the lamp base in Figure 12.
In lieu of this, we take inspiration from the construction of Čech complex in Topological Data Analysis. At a very high level, given a radius $r$ and two points, Čech complex forms the connectivity between the points by checking if the set of $r$-radius balls centered at each point intersect. If two $r$-radius balls have non-empty intersection, then they are connected, and vice versa. In our case of triangle meshes, the basic element is a triangle. If we perceive each triangle as a densely sampled set of points over the entire triangle, then the union of all $r$-radius balls for these points forms an offset surface from the triangle (see inset). Checking connectivity between triangles can then be done the same way as Čech complex by checking whether two offset surfaces have intersection. Intuitively, the construction leads to a straightforward implementation by computing triangle-to-triangle distances. Specifically, we first compute the closest points between two triangles, accelerated with a KD-tree. If the distance between their closest points is below a small threshold and the two triangles are not from the same connected component, we form a virtual edge between the two triangles. Since there are many possible vertex pairs to connect two triangles, we select the pair closest to the actual closest points (see the inset). Because our method additionally checks connected components, our result is not a Čech complex, but it shares a similar spirit in the construction.
Figure 12. Given a lamp mesh representing a single connected rigid object (first), the raw data, though, may contain several disconnected pieces from the modeling process (second). Using vertex distance (third) to infer the connectivity still leads to 4 disconnected components indicated by different colors [Garland and Heckbert, 1997]. In contrast, using our triangle distance (fourth) leads to a single component. We visualize both physical and virtual edges with thick black lines in the third and fourth images.
In Figure 12, we demonstrate that our triangle distance measure is more effective than the vertex distance proposed by Garland and Heckbert [1997]. By allowing merging of disconnected components, our method yields better simplification results (see Figure 13).
Figure 13. We simplify an input mesh (gray) with our method that computes virtual edges based on the triangle distances (blue). Our method can merge initially disconnected parts and lead to a more compact simplified mesh compared to the method by Garland and Heckbert [1997] which constructs virtual edges with vertex distances (green) and the method that is solely based on physical edges (red).
5.2. Vertex Pair Collapses
After constructing an edge set $E$ with physical and virtual edges, combining it with the input vertex and triangle lists forms a simplicial 2-complex $(V,E,F)$. Understanding the fact that our input is a simplicial 2-complex is important for choosing compatible data structures for implementing our method. But for brevity, we refer readers to section 8 about the implementation details and only touch the high-level idea about the simplicial complex.
We use greedy edge collapses to simplify the connectivity of a simplicial complex $(V,E,F)$. Collapsing an edge $ij\in E$ is a process of unifying a vertex pair $i,j$ to a new vertex $i^{\prime}$ and updating the edges and the faces connecting to them. More precisely, all the edges in $E$ and faces in $F$ that refer to vertex $i,j$ will be replaced by referring to vertex $i^{\prime}$. This process may introduce degenerated edges and faces
that connect to the same vertex $i^{\prime}$ more than once, or duplicated simplicies that connect to the same set of vertices (see inset). Thus, when updating $E$ and $F$, we remove the degenerated and duplicated edges and faces.
6. Error Metric
A key element in mesh simplification is to define an error metric that prioritizes the sequence of edge collapses. As pointed out in subsection 4.1, solely using the edge quadric error [Garland and Heckbert, 1997; Trettner and Kobbelt, 2020] may suffer from deleting large planar components.
Figure 14. Simplifying the input mesh with the edge quadric error may lead to deleting geometrically significant parts because they lead to nearly zero quadric error (green). Our method overcomes such issues by penalizing collapses that will introduce significant area change (blue).
Even a combination of the off-the-shelf tools, specifically [L98+T20] and [G97+L98+T20] mentioned in subsection 4.1, still suffer from other issues, such as overly smoothing the shape (see Figure 15). This motivates us to study a better metric to handle non-manifold meshes with multiple connected components.
Our key observation comes from studying the over smoothing artifact of our invented baseline and notices that the source error comes from how the area quadric is accumulated. When one contracts a virtual edge, it leads to gluing boundary edges (red edges in the inset) into interior edges (green edges in the inset). In such cases, the usual implementation of summing up quadrics (suggested in [Garland and Heckbert, 1997]) leads to area quadrics being applied to interior edges later on in the decimation. This behavior results in incorrectly high costs for removing interior edges. Collapsing an interior edge often leads to some area gain on one side of the edge and loss on the other. Although they should cancel out each other and result in a small area change, area quadrics do not differentiate between the two (see subsection 3.3) and would simply square the area changes on both sides and sum up the results. This results in a high cost when decimating long interior edges and leads to uniform decimation that does not preserve sharp features.
To overcome the issue of oversmoothing, we propose a small, but effective, change in the quadric error metric accumulation. Specifically, our quadric error metric follows the same combination as our proposed baseline [L98+T20] , combining the edge quadric [Garland and Heckbert, 1997] with the area quadric [Lindstrom and Turk, 1998]. But in contrast to our proposed baseline, we propose a different strategy to accumulate the quadrics. We accumulate the edge quadric with the usual summation as proposed in [Garland and Heckbert, 1997], a.k.a. the memory implementation, but we do not accumulate the area quadric term, a.k.a. the memoryless implementation. This small adjustment significantly improves the preservation of sharp features (see Figure 15). Our method retains the advantages of the quadric error decimation while avoiding the incorrect area quadric accumulation, as highlighted in the earlier observation.
Figure 15. Our novel quadric accumulation, the memory edge quadric with the memoryless area quadric, leads to a significantly better result in preserving sharp features compared to the straightforward combination of quadric [Garland and Heckbert, 1997] and area quadric [Lindstrom and Turk, 1998]. To illustrate the difference, we compare our method against such a combination, augmented with a tiny bit of probabilistic quadric [Trettner and Kobbelt, 2020]) for numerical robustness and extended it with the data structure of [Popovic and Hoppe, 1997] to handle non-manifolds. Given a soup of triangles (left), our method (right) leads to a better simplification result than the baseline (middle).
An interesting side observation is that we can achieve comparable results if we replace the area quadric with the boundary quadric presented in [Garland and Heckbert, 1998]. The reason for this is that the combination of the boundary quadric and the triangle quadric is equal to the squared height of the triangle incident to the boundary edge. When multiplied by the squared length of the boundary edge, this results in the squared triangle area, the same geometric quantity measured by the area quadric. This connection leads to the insight that replacing the area quadric with the boundary quadric is almost equivalent, except for a minor difference in the weighting between each term. Thus, we choose the area quadric because it is a more direct measure of the quantity (area change) we want to capture.
7. Texture Transfer with Successive Mapping
Most meshes come with textures. Preserving textures during simplification is essential for practical applications. Based on our observation that preserving UVs is not a robust approach to preserving textures (see subsection 4.2), we argue that a better way is to focus on transferring colors that existed on the surface, ignoring any UV boundary and UV islands. To achieve this, we maintain a surface correspondence map from the simplified model to the input model with successive mappings (e.g., [Liu et al., 2020]). With such a map, we can either directly render the simplified model with textures using the mesh color renderer [Yuksel et al., 2010] or regenerate a new texture image by copying the texture data from the original mesh with an entirely new UV mapping. We show that our successive mapping avoids texture bleeding in Figure 9 and also eliminates the resolution lower bound of UV-preserving simplification in Figure 11.
7.1. Successive Projection
Given an input mesh $\mathcal{M}^{0}=(V,E,F)$ and its coarsened counterpart $\mathcal{M}^{c}=(V_{c},E_{c},F_{c})$, our goal is to transfer surface attributes (e.g., texture colors) from $\mathcal{M}^{0}$ to $\mathcal{M}^{c}$. Without loss of generality, we consider each texel as a point sample defined on the surface. Transferring attributes amounts to compute a map $T:\mathcal{M}^{c}\rightarrow\mathcal{M}^{0}$ that transfers each point $p\in\mathcal{M}^{c}$ on the coarsened mesh to a corresponding point $T(p)\in\mathcal{M}$ on the input mesh. Then, for any point $p$ on $\mathcal{M}^{c}$, we can directly copy the attribute value from $T(p)\in\mathcal{M}^{0}$. With such information, we can generate a new texture map with attribute values copied from the input mesh $\mathcal{M}^{0}$.
We use the mesh color texture [Yuksel, 2017] for storing the resulting texture data after simplification. Specifically, given an input mesh with textures $\mathcal{M}^{0}$ (Figure 16 A), we compute the simplified mesh $\mathcal{M}^{c}$ with the mapping $T$ between the two. To bake the mesh color texture, we sample points on each triangle in $\mathcal{M}^{c}$ (Figure 16 C) and then use the mapping to obtain their corresponding points $T(p)$ on $\mathcal{M}^{0}$ (Figure 16 D). This correspondence allows us to pull back texture information from the input mesh $\mathcal{M}^{0}$ to color these sample points $\{p\}$ on $\mathcal{M}^{c}$ (Figure 16 E). Then we can either use the mesh colors [Yuksel et al., 2010] to visualize the result or bake them into a texture map (Figure 16 F).
Figure 16. Given an input mesh (A), we simplify the mesh (B) while computing a successive map from the simplified mesh to the input. With such information, we can sample points on the simplified mesh (C), correspondences to the pixel location in the texture image, and transfer those points to the input with the correspondence map (D). Then we can simply look up the color information on the input (E) and generate a new texture image after simplification (F).
Note that we can use any texture mapping technique to bake the texture map in Figure 16 as long as the mapping is injective: each texel in the texture image is either unused or corresponds to precisely a single point on the surface (see Figure 17). In our implementation, we simply use the trivial per-triangle flattening recommended in [Yuksel, 2017] due to its simplicity.
Figure 17. Our method for textured mesh simplification is applicable to any texture mapping method, as long as they can parameterize a simplicial complex. We show that both the mesh color textures [Yuksel, 2017] and the Blender’s smart-UV can be used to regenerate UV-coordinates.
Ideally, we want the mapping $T$ from $\mathcal{M}^{c}$ to $\mathcal{M}^{0}$ to be bijective with low distortion. However, since our method allows topological changes (e.g., deleting a component), bijectivity is ill-defined. Furthermore, because we make no assumption on the input, the original mesh may contain non-manifold vertices/edges and prevent us from using off-the-shelf manifold parameterization techniques to minimize distortion.
We thus compute the map $T$ by successive closest point projection within the edge one-ring. Specifically, when decimating $\mathcal{M}^{0}$, we construct a progressive simplicial complex [Popovic and Hoppe, 1997] so that we can go back and forth between the decimation sequence $\{\mathcal{M}^{0},\mathcal{M}^{1},\cdots,\mathcal{M}^{c}\}$. After decimation, we sample mesh color points on the coarsened mesh $\mathcal{M}^{c}$, use vertex split to obtain $\mathcal{M}^{c-1}$, and project the mesh color samples within the vertex one-ring of $\mathcal{M}^{c}$ to the edge one-ring of $\mathcal{M}^{c-1}$ (see inset). When we move on to subsequent projections to e.g. $\mathcal{M}^{c-2}$, a key detail is to reuse the point locations sampled from $\mathcal{M}^{c}$, but use the vertex split from $\mathcal{M}^{c-1}$ to $\mathcal{M}^{c-2}$ to identify the relevant edge one-ring on $\mathcal{M}^{c-2}$ that should be involved in the projection step. And we continue the process of projecting within one-rings until we reach $\mathcal{M}^{0}$.
Reusing point locations on $\mathcal{M}^{c}$ for closest point queries, but using successive maps $\mathcal{M}^{c}\rightarrow\mathcal{M}^{c-1}\rightarrow\cdots\rightarrow\mathcal{M}^{0}$ to identify relevant edge one-rings is crucial to our results. This ensures that the sampled surface attributes are geometrically closer to the simplified mesh $\mathcal{M}^{c}$ and avoids accumulating distortions due to the composition of successive projections. Our successive projection encourages (but does not guarantee) projecting each point to the part of the mesh it comes from. This is different from directly projecting points from $\mathcal{M}^{c}$ to the closest points on $\mathcal{M}^{0}$ which may project a point to an unrelated region, such as the wrong side of a thin feature (see Figure 18).
Figure 18. On a thin shell cloth mesh with different colors (blue, yellow) on different sides, naive closest point projection between the simplified and the input meshes may lead to projecting points onto the wrong side (middle). In contrast, our successive closest point projection encourages the projection to lie on the side, leading to less projection error.
8. Implementation
In addition to the proposed improvements over the edge topology, the error metric, and the texture-preserving simplification part, reproducing our method requires changes in the implementation (e.g., data structures) to properly handle non-manifold meshes. Here, we focus on discussing the differences in implementation compared to the standard approach, such as in [Garland and Heckbert, 1997], and omit the aspects that remain the same (e.g. using a priority queue to prioritize edge costs, employing a linear solve to determine the optimal vertex positions, etc.).
Simplicial Complex Data Structure
Many off-the-shelf mesh simplification methods are implemented using the half-edge data structure [Muller and Preparata, 1978]. However, such a data structure is primarily designed to handle manifold meshes with implicit assumptions e.g. each edge is shared by one or two faces, making it inapplicable to defective meshes.
As mentioned in section 5, our input belong to a simplicial 2-complex. Several data structures for manipulating simplicial complexes have been proposed to handle general simplices (e.g., high dimensional simplices) and reduce the memory footprint [Floriani and Hui, 2005]. However, in our case, these more generic data structures are unnecessary. We only perform edge collapses on simplicial 2-complexes. Thus, instead of using general-purpose data structures, we use a list of lists (similar to [Popovic and Hoppe, 1997]) to represent the simplicial complex. Specifically, we construct three lists to store the star information for each (1) vertex to its one-ring edges, (2) vertex to its one-ring faces, and (3) edge to its one-ring faces. For each edge collapse, we use these lists to gather neighboring simplicies and update their connectivity correspondingly.
Validity Checks
A widely adopted practice in implementing mesh simplification is to include validity checks before each edge collapse to maintain its topology and mesh quality, e.g. App.C in [Liu et al., 2020]. However, since the input mesh to our method may already contain topological defects and we aim to modify the topology (e.g., merging disjoint components) for better simplification results, we skip all validity checks. But if the input is known to be a manifold mesh and preserving topology after simplification is desired, our method can still be augmented with validity checks to reject edges that violate them.
9. Results
We enjoy the efficiency of being an edge collapse method that exhibits $O(n\log n)$ time complexity in the worst case, but nearly linear time complexity for moderate mesh resolutions. In the inset, we report the total runtime decomposed into virtual edge construction section 5, edge collapses section 6, and the texture transfer section 7 parts. The runtime is evaluated at the duck mesh (bottom right of the inset) at different resolutions with around 56,000 texture sample points. We evaluate our method on a MacBook with the M1 processor.
Qualitatively, we demonstrate that our method for constructing virtual edges leads to improvements over the previous method by Garland and Heckbert [1997] in Figure 13 and Figure 7. Combined with our modified quadric error metric, our method prioritizes merging disconnected components and avoids deleting components with large surface area (see Figure 14). Our method for textured mesh simplification avoids texture bleeding that many UV-preserving simplification methods suffer (see Figure 9), even when the mesh is non-manifold and has multiple components (see Figure 1).
Table 1. We decimate the meshes of the Thingi10K dataset [Zhou and Jacobson, 2016] down to 0.1%/1%/10% of its original resolution and report the geometric errors (Hausdorff distance, mean squared Chamfer distance) averaging across the dataset. The meshes are normalized to have unit diagonal bounding boxes to avoid bias towards big meshes.
Methods Hausdorff×$\times 10^{-1}$ Chamfer×$\times 10^{-1}$
[Garland and Heckbert, 1998] 1.05/0.57/0.13 2.36/1.37/0.46
[Trettner and Kobbelt, 2020] 1.05/0.58/0.14 2.36/1.36/0.39
Ours 0.84/0.44/0.10 2.06/1.11/0.29
Figure 19. We present a subset of textured mesh simplification results from [Garland and Heckbert, 1998] and our method obtained from the Real-World Textured Things dataset [Maggiordomo et al., 2020]. We can observe that a majority of textured meshes (the first column) in the wild have multiple texture islands in the UV map (the second column). Using the existing UV-preserving simplification method (third column) suffers from texture bleeding (seeing the background color of the texture image on the simplified mesh) and significantly larger geometric distortion because the algorithm is striking a balance between preserving UV and the geometry simultaneously. In contrast, our method decouples the UV and geometry preservation. Our geometric decimation part solely focuses on geometry preservation. We transfer texture to the simplified mesh with a successive mapping approach that guarantees no texture bleeding appears.
9.1. Quantitative Geometric Evaluations
We quantitatively evaluate the robustness of our method and show that our method can successfully simplify 100% of the models from the Thingi10k dataset [Zhou and Jacobson, 2016] down to 0.1% of its original resolution provided that the input mesh has more than 1000 faces (see Figure 20). In addition to robustness, our method produces results with a smaller geometric error, measured using the Hausdorff and the mean squared Chamfer distances (see Table 1).
Figure 20. Our method can robustly decimation all 10 thousand meshes from [Zhou and Jacobson, 2016] down to 1% of their original resolution. Here we display a randomly selected subset of the simplified meshes.
9.2. Quantitative Texture Evaluations
We also evaluate the quality of our texture preserving simplification. We compare our mesh simplification against [Garland and Heckbert, 1998] (implemented in [Cignoni et al., 2008]) on the Real-World Textured Things dataset [Maggiordomo et al., 2020]. Specifically, we decimate the $\approx$400 meshes in the dataset that contain a single texture image (though potentially with multiple texture islands) down to 0.1% of the input resolution. To avoid bias towards any specific render view, we measure the error using the average symmetric Chamfer distance on textures [Yuan et al., 2018], which measures the L2 norm of the color difference (color is normalized to [0,1]) between the closest spatial point pairs sampled on the surface. Our method results in 0.078 color error while the method by Garland and Heckbert [1998] leads to 0.201 error. In Figure 19, we show a subset of the decimation results from [Maggiordomo et al., 2020] to demonstrate the visual significance of the quantitative difference.
9.3. Perceptual User Studies
Besides all qualitative and quantitative evaluations, we perform a user study to gather perceptual opinions on mesh simplification results. Specifically, the participants were told to perform the user study in the context of a common online game scenario where the players are using low-end devices (e.g., mobile phones) to play online video games with other players. Due to the high performance requirement (e.g., frame rates, bandwidth), the 3D assets in the game have to satisfy a fixed budget (e.g., the total number of vertices in the scene). Under this situation, we asked participants to fill out an online survey (see inset) about their preferences in a variety of simplified 3D assets obtained from our method, the probabilistic quadric error simplification [Trettner and Kobbelt, 2020], and our invented baseline [G97+L98+T20] mentioned in section 4. The results from these methods are presented in an arbitrary order in the survey to avoid order bias. We invited 69 participants (51 male, 17 female, 1 non-binary individual, age ranging from 11-50, 75% graduate students and 25% working professionals including engineers and artists) to conduct the user study. Each of them was asked to select their favorite mesh simplification results out of the 10 user-created meshes presented in the paper (see Figure 22 for the collection) representing a wide range of meshes from organic shapes, 3D scenes, and man-made objects. From the total 690 responses (69$\times$10), 82.5% of them favored our results, 10.4% favored the results from [Trettner and Kobbelt, 2020], and 7.1% favored [G97+L98+T20]. The result demonstrates that our method leads to consistently better results across different meshes and individuals.
10. Discussion & Conclusion
We present a practical mesh simplification method that can achieve better results on human-created meshes that commonly appear in practice. We have also described an effective solution for copying the surface attributes without impacting the quality of the mesh simplification process. Our method pushes the robustness of mesh simplification to gracefully handle any triangle meshes, including non-manifold meshes and meshes with multiple connected components. This enables smoother performance by reducing the complexity of 3D models without compromising visual quality, enabling fast rendering and interactive simulation for real-time applications like online gaming, augmented reality, and virtual reality, especially on low-power devices such as mobile phones and VR headsets, where computational resources are limited.
Despite being robust, our method does not guarantee that the copied texture data comes from the out-most parts of the surface. Therefore, when simplifying meshes with multiple layers, the resulting texture can contain data from interior layers, which can be undesirable (see Figure 21).
Figure 21. Our texture is based on a localized closest point projection. When the decimation introduces topological changes (e.g. deleting components), our projection may pick up colors from different components, leading to noticeable texture distortion.
Future work can address this problem by considering external visibility information while computing the mapping from the final mesh to the input surface. Our research contributes to the field of robust geometry processing and underscores the need to further enhance the robustness of downstream geometric algorithms, particularly because the output of our method, like the input, may contain defects such as non-manifold structures.